There are 24 people. 9 have black hair, 4 have brown hair, 5 have blonde hair, and 6 have red hair. 3 people are then selected at random without replacement.

Solve the following probabilities:

(a) What is the probability that all 3 people selected have red hair.

(b) What is the probability that 2 people selected have the same color hair color, and the other person selected has a different hair color
Answer #1

Total = 24

Black hair = 9 , Brown Hair = 4 , Blonde Hair = 5 , Red Hair = 6

a) P(All 3 red ) = 6 C 3 / 24 C 3

= 20/2024

= 5/506

= 0.0099

b) P( 2 with same color ) = {P(2 black)*22 +P(2 brown)*22 + P(2 Red)*22 +P(2 blonde)*22} / 24 C 3

= 22(9 C 2 + 4 C 2  + 5 C 2 + 6 C 2 )/ 24 C 3

= 22(36+6+10+15)/2024

= 0.7283